OpenStudy (anonymous):

How do i solve 1/2 + (2/3 / 3/4) - (4/5 * 5/6) = I would like explanations and breakdown of how

OpenStudy (ujjwal):

Is it \[\frac{1}{2}+(\frac{\frac{2}{3}}{\frac{3}{4}})-(\frac{4}{5}*\frac{5}{6})\]

OpenStudy (anonymous):

yes

OpenStudy (ujjwal):

\[\frac{\frac{2}{3}}{\frac{3}{4}}=\frac{2}{3}*\frac{4}{3}\]can you do it now?

OpenStudy (anonymous):

how did u get it there

OpenStudy (ujjwal):

what?

OpenStudy (anonymous):

how did u get to that point what are the steps to solve it like dont you gotta change all denomenators to lcm before you do anything but the lcm would be 12 except 5 so it throws me off

OpenStudy (anonymous):

\[=1/2+(2/3 \div 3/4)-(4/5 \times 5/6)\] \[=1/2+(2/3 \times 4/3)-4/6\] \[=1/2+8/9-4/6\] \[=((1 \times 9)+(8 \times 2)-(4 \times 3))/18\] \[=13/18\]

OpenStudy (anonymous):

ty!!!!

OpenStudy (anonymous):

u r wlcm :)